# A fast, flexible simulation framework for Bayesian adaptive designs - the R package BATSS

### v1

@@ 3.6 Design specification - interim analyses and adaptations
The efficacy and futility stopping rules (for the intervention arms) are provided via user-defined functions in the eff.arm and fut.arm arguments, respectively (with input parameters provided in eff.arm.control and fut.arm.control). Efficacy and futility stopping rules of the trial at a given interim analysis are respectively indicated via the eff.trial and fut.trial arguments that take the target parameters results as input. This, for example, allows the trial to stop once one intervention arm has been found efficacious. Several example functions are provided in the BATSS vignette (Couturier et al., 2024). We defer their detailed explanation to the case studies (Section 4, and Appendices B and C).
@@ 3.7 BATSS ingredients
The batss.glm function calculates a number of “ingredients” or derived quantities at each interim analysis which can be used as inputs to the user-defined functions for the adaptations. These are summarised in Table 1 (full description given in Table A2 in Appendix A, and in the vignette (Couturier et al., 2024)).
@@ 3.8 Parallel and cluster computing
In BATSS, the number of simulated trials is provided by means of the argument R, either with a scalar denoting the number of trials (in that case seeds 1 to R will be used, e.g., 1 to 10000 for our example above) or with a vector of seeds to use to generate each trial’s data (in that case the the number of trials will equal the length of that vector).
@@
  • To split the computational burden between different computers and/or the CPUs of one or several clusters, the user can split the planned list of seeds into non-overlapping sets. Each set will then be assigned to an available resource (like a CPU on a cluster), either by hand or more efficiently via a script (like an sbatch on clusters using Simple Linux Utility for Resource Management (SLURM) as the job manager), by a call of the batss.glm function using the set of seeds of interest through the R argument and saving the results of each trial (setting argument extended $>$ 0). The batss.combine function, taking as input the location of the R data files containing the outputs corresponding to each set of seeds, will then be used to aggregate results. (Note that batss.combine checks if the sets of seed overlap or if the calls to batss.glm were different).
@@ 3.9 Output
The summary method (obtained by using the summary function on the output of batss.glm) first prints a summary of the design and inputs to the batss.glm function. Then, in its “short” variant (default), it indicates the probability of declaring efficacy or futility for each intervention arm, for at least one intervention arm, as well as for all intervention arms for each hypothesis of interest (i.e., under the null and/or alternative hypotheses if argument H0=TRUE, for example). Note that, as a diagnostic tool, the probability of simultaneously meeting both the efficacy and futility criteria is also provided if this highly undesirable situation occurred for any arm in any simulated trial. In its “long” variant (obtained using the summary function with full = TRUE option), the method additionally provides: a breakdown of when the efficacy or futility stopping occurred (early = at an interim analysis; last = at the maximum sample size); summaries of the sample sizes for each arm and overall (mean, standard deviation, median and 10th and 90th percentiles); and a summary of the different scenarios that occurred in terms of the decisions for each intervention arm (no decision, efficacy decision, futility decision, both efficacy and futility criteria met (this is only displayed for simulated designs where this undesirable behaviour has occurred)) and how frequently those different combinations occurred across the Monte Carlo trials (and what proportion of those had early decisions made where the trial stopped at an interim analysis).
@@
Plots of the posterior mean of the target parameter (i.e., treatment effect) estimates versus the sample size at the last look performed for that Monte Carlo trial, along with the decision (efficacy, futility, none, both), for each intervention arm, can be obtained using the plot function with the type = "estimates" option (demonstrated in Figure 2). Plots of the distributions of the sample sizes (across the Monte Carlo trials) can also be obtained using the type = "size" option in the plot function (demonstrated in Figure 3).
@@ 4 Case Study 2 - binary outcome
    $\frac{P(\beta_{k}>\delta_{r}|\ \mathbf{y},\mathbf{X})^{h}}{\sum^{K_{j}}_{k=1}P(\beta_{k}>\delta_{r}|\ \mathbf{y},\mathbf{X})^{h}}$ (4)
@@
  prob = rep(NA,g) #allocation probabilities
  # reference/control arm allocation - Equation 3
@@
  # targets/interventions (that haven’t been dropped) - Equation 4
@@
(The arguments that are boldfaced are BATSS ingredients, and those that are italicised are items from the respective user-provided ‘.control’ arguments in the batss.glm function. This same convention applies throughout this section.)
@@
  prob = abs(prob)/sum(abs(prob)) #normalise probabilities
@@
futility.arm.fun = function( posterior, b.fut){
@@
futility.arm.fun.inc = function( posterior, n,N,b.fut,p.fut){
@@
  #delta.eff needs to have NAs to not evaluate efficacy at interims
@@
The pairwise one-sided type I error was calculated as the proportion of simulations that falsely declared efficacy when the intervention was assumed to have the same ORR as the control (i.e., log OR = 0). The (pairwise) power was calculated as the proportion of simulations that declared efficacy for an intervention when it was assumed to have a large effect (i.e., $>20\%$ absolute improvement in the ORR compared to the control; as in the “alternative” scenarios).
@@
Here we have focused on providing the user-defined functions for performing the adaptations. The full description of the batss.glm() call for simulating the “Alternative” scenario for Examples 2.1 and 2.2 can be found in Appendix C, following the logic described above.
@@ 4.1 Results
Table 3 shows the proportion of simulations that declared each intervention arm to be effective (at the maximum sample size/end of the trial), the proportion of simulations that each intervention arm was stopped early to declare futility, and the average sample size per arm and overall sample size, for each scenario for Case Study 2.1. A summary of the target parameter estimates against the sample size achieved for a simulated trial is provided in Figure 4. The distribution of sample sizes for each treatment arm can also be examined in Figure 5 for the “Alternative” scenario (provided for illustrative purposes).
@@
Table 3: Pairwise type I error, power, proportion of early stopping of intervention arms for futility, and sample sizes for the binary outcome case study with a fixed futility boundary value (Case Study 2.1)
A (Control) B C D E F Overall
Scenario Probability of declaring a treatment effective
Null NA 0.044 0.043 0.046 0.044 0.045 0.154*
Alternative NA 0.046 0.041 0.211 0.900 0.902 0.982*
Scenario Probability of stopping an intervention arm early for futility
Null NA 0.588 0.593 0.597 0.590 0.588 -
Alternative NA 0.508 0.509 0.279 0.033 0.028 -
Scenario Average (SD) sample size
Null 49.1 (17.3) 28.4 (18.0) 27.9 (17.7) 27.6 (17.7) 28.1 (17.9) 28.1(18.0) 189.2 (48.7)
Alternative 49.9 (9.5) 20.6 (9.8) 20.3 (9.7) 29.0 (12.2) 47.7 (10.6) 48.0 (10.4) 215.5 (8.4)
  • *The proportion of simulations that declared at least one intervention arm to be effective
Figure 4: Scatterplot of target parameter estimates (posterior mean) versus sample size for each intervention arm (separate plots) at the time of stopping the arm, along with the decision at the time of stopping for the “Alternative” scenario for Case Study 2.1 across the 10’000 simulated trials. The timing of the stop is also provided (“Early stopping” = stopping at an interim analysis).
Figure 5: Box plot showing the distribution of overall sample size and sample size for each treatment arm across the 10’000 simulated trials for the “Alternative” scenario for Case Study 2.1.
@@
Table 4: Pairwise type I error, power, proportion of early stopping of intervention arms for futility, and sample sizes for the binary outcome case study with an increasing futility boundary value (Case Study 2.2)
A (Control) B C D E F Overall
Scenario Probability of declaring a treatment effective
Null NA 0.030 0.031 0.0300 0.029 0.029 0.097*
Alternative NA 0.040 0.037 0.186 0.864 0.872 0.968*
Scenario Probability of stopping an intervention arm early for futility
Null NA 0.945 0.949 0.947 0.950 0.949 -
Alternative NA 0.931 0.935 0.750 0.116 0.106 -
Scenario Average (SD) sample size
Null 39.7 (14.8) 23.0 (13.6) 22.7 (13.5) 22.4 (13.5) 22.9 (13.6) 22.9 (13.7) 153.7 (45.9)
Alternative 53.9 (10.1) 18.7 (8.7) 18.5 (8.6) 26.2 (12.0) 48.3 (12.1) 48.8 (11.9) 214.4 (12.5)
  • *The proportion of simulations that declared at least one intervention arm to be effective
@@ 5 Discussion
Here we have presented an R package, BATSS (Couturier et al., 2024), which provides a modular, flexible, structured framework to simulate Bayesian adaptive trials. BATSS is an open source, script-based package that is extendable, and rich with options for adaptations for simulating Bayesian adaptive clinical trials. It allows for easy calculation of various trial operating characteristics and visualisation of some key metrics. We have provided an overview of the workflow and arguments required for the batss.glm function which can be used to simulate Bayesian adaptive trials whose outcome may be modelled by a GLM.
@@ 5.1 Limitations and Future Research
The BATSS package does not currently support platform trial designs where treatment arms may be added or those implementing multi-factorial designs (with multiple domains of treatment). This is an area for future research. We also plan to extend the package to support hierarchical models so that designs that implement information borrowing across subgroups can be implemented (e.g., basket trials). Hosseini et al. (2024) provide evidence that this is feasible and fast using INLA, and may provide more computational savings compared to MCMC for these models.
For simplicity, this version of BATSS does not simulate the enrollment times of patients or account for follow-up duration (leading to delays in data availability) or loss to follow-up. This is similar to most existing software for simulating Bayesian adaptive designs. The batss.glm function assumes that analyses are conducted based on the number of participants completing follow-up (which, here, is also the same as the number recruited). However, it is highly likely that there will be more participants recruited with incomplete follow-up at the time of an interim analysis, and so the sample sizes across the simulated trials will be slightly underestimated. This version of BATSS is appropriate for slow recruiting trials and/or those that have endpoints that are observed over a short period (and no loss-to-follow-up). We are currently extending BATSS to explicitly simulate accrual, drop-outs and account for the follow-up duration length (and therefore have participants with incomplete follow-up in the trial at interim analyses), and the next release of BATSS will contain these features.

### v2

@@ 3.6 Design specification - interim analyses and adaptations
The efficacy and futility stopping rules (for the intervention arms) are provided via user-defined functions in the eff.arm and fut.arm arguments, respectively (with input parameters provided in eff.arm.control and fut.arm.control). Efficacy and futility stopping rules of the trial at a given interim analysis are respectively indicated via the eff.trial and fut.trial arguments that take the target parameters results as input. This, for example, allows the trial to stop once one intervention arm has been found efficacious. Several example functions are provided in the BATSS vignette (Couturier et al., 2025). We defer their detailed explanation to the case studies (Section 4, and Appendices B and C).
@@ 3.7 BATSS ingredients
The batss.glm function calculates a number of “ingredients” or derived quantities at each interim analysis which can be used as inputs to the user-defined functions for the adaptations. These are summarised in Table 1 (full description given in Table A2 in Appendix A, and in the vignette (Couturier et al., 2025)).
@@ 3.8 Parallel and cluster computing
In BATSS, the number of simulated trials is provided by means of the argument R, either with a scalar denoting the number of trials (in that case seeds 1 to R will be used, e.g., 1 to 10,000 for our example above) or with a vector of seeds to use to generate each trial’s data (in that case the number of trials will equal the length of that vector).
@@
  • To split the computational burden between different computers and/or the CPUs of one or several clusters, the user can split the planned list of seeds into non-overlapping sets. Each set will then be assigned to an available resource (like a CPU on a cluster), either by hand or more efficiently via a script (like an sbatch on clusters using Simple Linux Utility for Resource Management (SLURM) as the job manager), by a call of the batss.glm function using the set of seeds of interest through the R argument and saving the results of each trial (setting argument extended >> 0). The batss.combine function, taking as input the location of the R data files containing the outputs corresponding to each set of seeds, will then be used to aggregate results. (Note that batss.combine checks if the sets of seed overlap or if the calls to batss.glm were different).
@@ 3.9 Output
The summary method (obtained by using the summary function on the output of batss.glm) first prints a summary of the design and inputs to the batss.glm function. Then, in its “short” variant (obtained by opting for extended = FALSE), it indicates the probability of declaring efficacy or futility for each intervention arm, for at least one intervention arm, as well as for all intervention arms for each hypothesis of interest (i.e., under the null and/or alternative hypotheses if argument H0=TRUE, for example). Note that, as a diagnostic tool, the probability of simultaneously meeting both the efficacy and futility criteria is also provided if this highly undesirable situation occurred for any arm in any simulated trial. In its “long” variant (obtained using the summary function with extended = TRUE option), the method additionally provides: a breakdown of when the efficacy or futility stopping occurred (early = at an interim analysis; last = at the maximum sample size); summaries of the sample sizes for each arm and overall (mean, standard deviation, median and 10th and 90th percentiles); and a summary of the different scenarios that occurred in terms of the decisions for each intervention arm (no decision, efficacy decision, futility decision, both efficacy and futility criteria met (this is only displayed for simulated designs where this undesirable behaviour has occurred)) and how frequently those different combinations occurred across the Monte Carlo trials (and what proportion of those had early decisions made where the trial stopped at an interim analysis).
@@
Plots of the posterior mean of the target parameter (i.e., treatment effect) estimates versus the sample size at the last look performed for that Monte Carlo trial, along with the decision (efficacy, futility, none, both), for each intervention arm, can be obtained using the plot function with the which = 4 option (demonstrated in Figure 4). Plots of the distributions of the sample sizes per group (across the Monte Carlo trials) can also be obtained using the which = 3 option in the plot function (demonstrated in Figure 3), as well as using the which = 1 option.
@@ 4 Case Study 2 - binary outcome
    $\frac{P(\beta_{k}>\delta_{r}|\ \mathbf{y},\mathbf{X})^{h}}{\sum^{K_{j}}_{l=1}P(\beta_{l}>\delta_{r}|\ \mathbf{y},\mathbf{X})^{h}}$ (4)
@@
# vector of allocation probabilities:
prob = rep(NA,g)
# reference/control arm allocation - Equation 3:
@@
# targets/interventions (that haven’t been dropped) - Equation 4:
@@
where the arguments in orange are BATSS ingredients (Table 1), and those in red are items from the respective user-provided ‘.control’ arguments in the batss.glm function. This same convention applies throughout this section.
@@
prob = abs(prob)/sum(abs(prob)) # (normalise probabilities)
@@
futility1.arm.fun = function(posterior, b.fut){
@@
futility2.arm.fun = function(posterior, n, N, b.fut, p.fut){
@@
# where delta.eff needs to have NAs to not evaluate efficacy at interims
@@
The pairwise one-sided type I error was calculated as the proportion of simulations that falsely declared efficacy when the intervention was assumed to have the same ORR as the control (i.e., log OR = 0). The (pairwise) power was calculated as the proportion of simulations that declared efficacy for an intervention when it was assumed to have a large effect (i.e., $20\%$ absolute improvement in the ORR compared to the control; as in the “alternative” scenario).
@@
Here we have focused on providing the user-defined functions for performing the adaptations. The full description of the batss.glm() call for simulating the “Alternative” scenario for Case Studies 2.1 and 2.2 can be found in Appendix C, following the logic described above.
@@ 4.1 Results
Table 3 shows the proportion of simulations that declared each intervention arm to be effective (at the maximum sample size/end of the trial), the proportion of simulations that each intervention arm was stopped early to declare futility, and the average sample size per arm and overall sample size, for each scenario for Case Study 2.1. Figure 5 shows the probability of stopping at a given look and confirms that early stopping – only occurring when all arms are found futile – is rare. The distribution of sample sizes for each treatment arm can also be examined in Figure 6 for the “Alternative” scenario. A summary of the target parameter estimates against the sample size achieved for a simulated trial is provided in Figure 7.
@@
Conclusion Sample size
Group Efficacy(1) Futility(2) Average(3) Stand. Dev.(4)
Under the global null hypothesis:
Control (Arm A) - - 49.13 17.31
Arm B 0.044 0.604 28.36 17.97
Arm C 0.044 0.609 27.91 17.81
Arm D 0.046 0.611 27.62 17.68
Arm E 0.046 0.608 28.13 17.93
Arm F 0.045 0.605 28.09 17.98
At least 1 0.156 0.890 - -
All 0.001 0.313 189.24 48.61
Under the alternative hypothesis:
Control (Arm A) - - 49.89 9.46
Arm B 0.046 0.516 20.59 9.76
Arm C 0.040 0.518 20.32 9.65
Arm D 0.211 0.284 28.98 12.20
Arm E 0.899 0.033 47.72 10.62
Arm F 0.901 0.028 48.03 10.44
At least 1 0.982 0.723 - -
All 0.006 0.004 215.51 8.27
Table 3: (1) Probability of concluding a treatment arm, at least one treatment arm, or all treatment arms (arms B to F) to be effective at the end of the trial, (2) probability of stopping early a treatment arm, at least one treatment arm, or all treatment arms for futility (arms B to F), (3)(4) expected sample size per arm (A to F) and in total, and corresponding standard deviation, under the null and alternative hypotheses (refer to Table 2) in Case study 2.1
Figure 5: Barplot of the probability of stopping (y-axis) at each look (x-axis) as observed across the 10,000 simulated trials for the “Alternative” scenario for Case Study 2.1. The vertical blue dashed line shows the estimate of the expected sample size (ESS). As the trial only stops if all arms are dropped for futility, the expected sample size almost equals 216, the maximum sample size.
Figure 6: Violin plot (and boxplot) of the simulated sample sizes per group and corresponding expected sample sizes for the “Alternative” scenario for Case Study 2.1 across the 10,000 simulated trials. The sample size per group of each trial are displayed. For target parameters, these points are colour-coded and shaped by the decision and timing when the trial stopped. We can note that the ineffective arms ‘B’ and ‘C’ are are often dropped early for futility and that more patients are allocated to effective arms ‘E’ and ‘F’.
Figure 7: Scatterplot of target parameter log odds ratio estimates (posterior mean) versus sample size for each intervention arm (separate plots) at the time of stopping the arm, along with the decision at the time of stopping for the “Alternative” scenario for Case Study 2.1 across the 10,000 simulated trials. The timing of the stop is also provided (“Early stopping” = stopping at an interim analysis). We can note that log odds ratio estimates are typically negative when dropping arms early for futility.
@@
Conclusion Sample size
Group Efficacy(1) Futility(2) Average(3) Stand. Dev.(4)
Under the global null hypothesis:
Control (Arm A) - - 39.75 14.79
Arm B 0.030 0.965 22.98 13.52
Arm C 0.031 0.964 22.71 13.56
Arm D 0.030 0.965 22.46 13.49
Arm E 0.030 0.966 22.87 13.59
Arm F 0.029 0.965 22.94 13.78
At least 1 0.097 0.999 - -
All 0.000 0.883 153.71 45.82
Under the alternative hypothesis:
Control (Arm A) - - 53.89 10.09
Arm B 0.041 0.956 18.66 8.65
Arm C 0.037 0.960 18.47 8.54
Arm D 0.189 0.799 26.27 12.02
Arm E 0.863 0.126 48.32 12.13
Arm F 0.874 0.116 48.86 11.87
At least 1 0.968 0.993 - -
All 0.006 0.027 214.46 12.20
Table 4: (1) Probability of concluding a treatment arm, at least one treatment arm, or all treatment arms (arms B to F) are effective at the end of the trial, (2) probability of stopping early a treatment arm, at least one treatment arm, or all treatment arms for futility (arms B to F), (3)(4) expected sample size per arm (A to F) and in total, and corresponding standard deviation, under the null and alternative hypotheses (refer to Table 2) in Case study 2.2
@@ 5 Discussion
Here we have presented an R package, BATSS (Couturier et al., 2025), which provides a modular, flexible, structured framework to simulate Bayesian adaptive trials. BATSS is an open source, script-based package that is extendable, and rich with options for adaptations for simulating Bayesian adaptive clinical trials. It allows for easy calculation of various trial operating characteristics and visualisation of some key metrics. We have provided an overview of the workflow and arguments required for the batss.glm function which can be used to simulate Bayesian adaptive trials whose outcome may be modelled by a GLM.
@@ 5.1 Limitations and Future Research
The BATSS package does not currently support platform trial designs where treatment arms may be added or those implementing multi-factorial designs (with multiple domains of treatment). This is an area for future research. We also plan to extend the package to
  • support hierarchical models so that designs that implement information borrowing across subgroups can be implemented (e.g., basket trials). Hosseini et al. (2024) provide evidence that this is feasible and fast using INLA, and may provide more computational savings compared to MCMC for these models.
  • consider the enrollment times of patients or account for follow-up duration (leading to delays in data availability) or loss to follow-up and drop-out. Instead of assuming that analyses are conducted based on the number of participants completing follow-up (which, here, is also the same as the number recruited), a related extension would be to take into account the fact that some of the participants recruited have incomplete follow-up at the time of an interim analysis. Such features, typically not included in existing software for simulating Bayesian adaptive designs, would allow estimating further operating characteristics, mentioned by the guidelines described in Section 1.1, like, for example, the expected, minimum and maximum study times and number of recruited participants ($F^{19}_{2}$, $I^{20}_{1}$). We are currently extending BATSS to explicitly simulate accrual, drop-outs and account for the follow-up duration length (and therefore have participants with incomplete follow-up in the trial at interim analyses), and the next release of BATSS will contain these features. This is necessary in particular for survival/time-to-event endpoints.
  • propose point and interval estimators for the target parameters. Bias and mean squared error of point estimators as well as coverage of interval estimators are typically of interest for regulators ($F^{11}_{2}$, $I^{11}_{1}$, $I^{21}_{1}$. Whilst raw trial data for each simulated data set are available to users (by specifying extended = 2 in batss.glm) to define the properties of their planned estimators (see, e.g., Robertson et al. (2023a), Robertson et al. (2023b), Robertson et al. (2025a), Robertson et al. (2025b)), allowing this task to be performed with the same modular approach as to define other aspects of the trial, would be useful.
This version of BATSS is appropriate for slow recruiting trials and/or those that have endpoints that are observed over a short period (and no loss-to-follow-up).